# Conceptualization, Methodology, Writing – original draft, Data curation, Software, Writing – review & editing

### v1

@@ 1 Introduction
Traffic oscillation, alternatively termed stop-and-go traffic, is a recurrent pattern of deceleration and acceleration typically observed in congested traffic scenarios. This phenomenon significantly impairs traffic efficiency, compromises safety, and diminishes energy efficiency (Bilbao-Ubillos, 2008; Pu et al., ; Tian et al., 2024). Extensive research has substantiated that a multitude of maneuvers can instigate traffic oscillations, including lane changing (Ahn and Cassidy, 2007; Mauch and J. Cassidy, 2002; Laval and Daganzo, 2006), merging, and diverging (Ahn et al., 2010; Cassidy and Rudjanakanoknad, 2005). The rapid progression of automated vehicles (AVs) technologies (Guanetti et al., 2018; Rajamani, 2011; Li et al., 2025b, a; Tian et al., 2025) introduces a new traffic flow system: mixed traffic comprising a heterogeneous mix of AVs and human-driven vehicles (HDVs). This mixed traffic system is expected to reshape traditional traffic flow dynamics in unparalleled ways and improve traffic throughput, safety as well as energy efficiency during traffic oscillations (Talebpour and Mahmassani 2016; Li et al. 2024a; Xu et al. 2022; Anis et al. 2025). Consequently, acquiring a profound understanding of the distinctive behaviors exhibited by AVs during traffic oscillations becomes imperative in order to model and analyze this new traffic flow system.
@@
[…]
@@ 2 Modeling the CF problem of AVs
where $a_{n+1}(t)$ is the acceleration of the AV $n+1$ at time $t$, $\Delta d_{n+1}(t)=p_{n}(t)-p_{n+1}(t)-d^{*}_{n+1}(t)$ is the spacing deviation of AV $n+1$ at time $t$, and $d^{*}_{n+1}(t)=l_{n+1}+\tau\dot{p}_{n+1}(t)$ is the desired spacing of AV $n+1$ at time $t$ wherein $l_{n+1}$ is a constant distance, $\tau$ is a constant time, and $\dot{p}_{n+1}(t)$ is the speed of AV $n+1$. $\Delta v_{n+1}(t)=\dot{p}_{n}(t)-\dot{p}_{n+1}(t)$ is the speed difference between AV $n+1$ and its predecessor at time $t$. $k_{d}$ and $k_{v}$ are the feedback gains on $\Delta d_{n+1}$ and $\Delta v_{n+1}$, respectively. It is imperative to note that $d^{*}_{n+1}(t)=l_{n+1}+\tau\dot{p}_{n+1}(t)$ is a generic form. For example, under the two most prevalent spacing policies—constant time gap policy (Naus et al., 2010; Kianfar et al., 2015) and constant distance policy (Swaroop and Hedrick, 1999; Naus et al., 2010; Li et al., 2024b)—$\tau$ equals to the pre-defined time gap and zero, respectively. Based on the trajectory decomposition shown in Fig. 2, we obtain:
@@
While Eq. (7) have also been applied to describe HDVs in the literature, they typically rely on strong assumptions about deterministic behavior and further linearization simplifications. To maintain a clear and rigorous scope, this study focuses on AVs.
@@
In practice, time delays could impact the response of AV systems. As this paper does not address connectivity issues, focusing on actuation delays is sufficient. We can introduce an actuation delay into our framework via the block diagram shown in Fig. 2, where the added delay block signifies an actuation delay of $\tau$ seconds. A time delay is known to be linear time-invariant, with a frequency response magnitude of 1 and a phase shift of $-\omega\tau$ (Franklin et al., 2002). To simplify the presentation of our framework, we do not incorporate the delay component in our analysis that follows. It is worth noting that, the steps of our analysis procedure remain consistent even when incorporating the delay, as it merely represents an additional linear component without introducing any further nonlinearity.
@@ 3.2.1 Control saturation
For the control saturation element (Saturation 1 in Fig. 2), let the input of it be denoted as $a_{n+1}(t)=B_{a}sin(\omega t)$, considering the acceleration limits as $[a_{min},a_{max}]$. The DF of the control saturation, $N_{a,sat}(B_{a})$, can be characterized into four cases corresponding to different limit activeness, depending on the value of $B_{a}$, given the known values of $a_{min}$ and $a_{max}$:
@@
Case (2): limit activeness by $a_{min}$ (i.e., $-a_{min}\leq B_{a}\leq a_{max}$), as illustrated in Fig. 3.2.1(b), where $\beta_{1}=sin^{-1}(\frac{-a_{min}}{B_{a}})$.
@@
$\dot{\tilde{v}}_{n+1}(t)=\begin{cases}B_{a}\sin(\omega t)&\text{if }0<\omega t\leq\beta_{2}\\ a_{\max}&\text{if }\beta_{2}<\omega t\leq\pi-\beta_{2}\\ B_{a}\sin(\omega t)&\text{if }\pi-\beta_{2}<\omega t\leq 2\pi\\ \end{cases}$ (17)
@@
The cases in this subsection only describe how a signal passes through the saturation element wihout considering the closed-loop effect. The closed-loop response is inherently much more complex than the open-loop case due to feedback interactions (illustrated in Fig. 2). A full characterization of the closed-loop behavior requires analyzing how signals propagate throughout the feedback loop. This is systematically addressed in Section 4, where we examine the overall system response.
@@ 3.2.2 State saturation
Similar, the DF of the state saturation (Saturation 2 in Fig. 2) with input $\tilde{v}_{n+1}(t)=B_{v}\sin(\omega t)$ and its output limited within $[\tilde{v}_{min},\tilde{v}_{max}]$ can be derived in the same manner:
@@
Case (2): limit activeness by $\tilde{v}_{min}$ (i.e., $-\tilde{v}_{min}\leq B_{v}\leq\tilde{v}_{max}$).
@@ 4.2.1 Systems with solely control saturation
$=\sqrt{(k_{1}R)^{2}+(\omega k_{2}R)^{2}}\sin\left(\omega t+\tan^{-1}\left(\frac{k_{2}\omega}{k_{1}}\right)\right)\quad\quad\text{if }1-\frac{k_{1}|N_{a,sat}(B_{a})|}{\omega^{2}}\geq 0$
@@
$=\sqrt{(k_{1}R)^{2}+(\omega k_{2}R)^{2}}\sin\left(\omega t+\tan^{-1}\left(\frac{k_{2}\omega}{k_{1}}\right)\right)\quad\quad\text{if }1-\frac{k_{1}|N_{a,sat}(B_{a})|}{\omega^{2}}<0$
@@ 4.2.2 Systems with solely state saturation
$=\sqrt{\left(\frac{-k_{1}R}{\omega}\right)^{2}+(k_{2}R)^{2}}\sin\left(\omega t-\tan^{-1}\left(\frac{k_{1}}{k_{2}\omega}\right)\right)\quad\quad\text{if }1-\frac{k_{1}|N_{v,sat}(B_{v})|}{\omega^{2}}\geq 0$
@@
$=\sqrt{\left(\frac{-k_{1}R}{\omega}\right)^{2}+(k_{2}R)^{2}}\sin\left(\omega t-\tan^{-1}\left(\frac{k_{1}}{k_{2}\omega}\right)\right)\quad\quad\text{if }1-\frac{k_{1}|N_{v,sat}(B_{v})|}{\omega^{2}}<0$
@@ 4.2.3 Systems with both control and state saturation
$\cdot\sin\left(\omega t+\phi_{a}-\tan^{-1}\left(\frac{-\omega k_{3}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}{\omega^{2}-k_{1}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}\right)\right)\quad\quad\text{if }1-\frac{k_{1}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}{\omega^{2}}\geq 0$
@@
$\cdot\sin\left(\omega t+\phi_{a}-\tan^{-1}\left(\frac{-\omega k_{3}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}{\omega^{2}-k_{1}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}\right)+\pi\right)\quad\quad\text{if }1-\frac{k_{1}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}{\omega^{2}}<0$
@@
$-\omega^{2}\tilde{p}_{n+1}(\omega)=k_{1}\tilde{p}_{n}(\omega)+j\omega k_{2}\tilde{p}_{n}(\omega)-k_{1}\tilde{p}_{n+1}(\omega)+j\omega k_{3}\tilde{p}_{n+1}(\omega)$ (52)
@@ 4.2.4 Solving for forced oscillation candidates
For the three types of systems analyzed above, given the values of $B_{a,s}$ or $B_{v,s}$, the variables $\phi_{a,s}$ or $\phi_{v,s}$ can be analytically determined with ease by equating the phase of both sides of equations Eqs. (35)-(36), Eqs. (41)-(42), or Eqs. (47)-(48) within the studied range of $\omega$. $B_{a,s}$ or $B_{v,s}$ can be deduced by equating the amplitude of the two sides of the aforementioned equations. However, the nonlinearity introduced by the DFs can render the analytical solution of these equations challenging. Several studies have proposed graphical methods to identify the roots of such equations (Vander Velde et al., 1968; Atherton, 2011; Li et al., 2012; Maraini and Nataraj, 2018). Nevertheless, implementing these graphical solutions for heterogeneous AVs can be cumbersome and repetitive. Fortunately, iterative strategies, such as the Brent-Dekker method (Dekker, 1969; Brent, 2013), offer numerical solutions. These strategies are largely dependent on an initial guess, hence, we sweep the initial guess across the range $[0,B_{ini,max}]$ to identify all possible solutions, where $B_{ini,max}$ represents a predetermined maximum value of the initial guess.
@@ 5.1 The incremental-input DF
where the first term is the predominant input which is one of the oscillation candidates acquired from the analysis in Section 4. It’s shifted on the time axis so that its first cycle starts at $t=0$, simplifying notation without affecting the analysis due to our interest in the phase relation between primary and incremental inputs. The second term represents the incremental input which can be viewed as a perturbation to the predominant input. In this context, $\epsilon\ll B_{s}$, indicating the perturbation’s magnitude is notably less than the predominant input. Additionally, $\theta$ represents the phase shift of the incremental input in relation to the predominant input, which can span any value in the interval $[0,2\pi]$.
@@ 5.2.1 Control saturation
For the control saturation element, let the predominant input of it be denoted as $a_{n+1}(t)=B_{a,s}sin(\omega t)$, where $B_{a,s}$ can be derived from Eqs. (35)-(36) or Eqs. (47)-(48). Considering the acceleration limits as $[a_{min},a_{max}]$. It has been shown that the DF of the control saturation, $N_{a,sat}(B_{a,s})$, can be characterized into four cases depending on different limit activeness. Analogously, the incremental-input DF, $N_{inc}(B_{a,s},\theta_{a})$, can also be classified into these four scenarios:
@@
Case (2): limit activeness by $a_{min}$ (i.e., $-a_{min}\leq B_{a,s}\leq a_{max}$).
@@ 5.2.2 State saturation
Similarly, consider the predominant input to the state saturation $\tilde{v}_{n+1}(t)=B_{v,s}\sin(\omega t)$, where $B_{v,s}$ can be obtained by solving Eqs. (41)-(42) or Eqs. (47)-(48), and the output of the saturation limited within $[\tilde{v}_{min},\tilde{v}_{max}]$. The incremental-input DF of the state saturation, $N_{v,inc}(B_{v,s},\theta_{v})$, can be obtained as follows:
@@
Case (2): limit activeness by $\tilde{v}_{min}$ (i.e., $-\tilde{v}_{min}\leq B_{v,s}\leq\tilde{v}_{max}$).
@@ 6.1 Simulation of system containing one saturation
The results demonstrate that upon reaching the saturation limits, the linear theoretical method, which is commonly employed in much of the current research on string stability, significantly overestimates the magnitude of the frequency response, $|F|$. In other words, an underestimation of the attenuation effect in traffic oscillation largely exists for linear analysis. On the other hand, the DF approximation introduced by Zhou et al., 2023 marginally overestimates the oscillation attenuation. Both the linear method and DF approximation considerably overestimate the phase of the frequency response, $\angle F$. Consequently, the response time (defined as the duration before the AV responds to a predecessor’s driving behavior change), which can be calculated by $-\frac{\angle F}{\omega}$, is underestimated.

### v3

@@ 1 Introduction
Traffic oscillation, alternatively termed stop-and-go traffic, is a recurrent pattern of deceleration and acceleration typically observed in congested traffic scenarios. This phenomenon significantly impairs traffic efficiency, compromises safety, and diminishes energy efficiency (Bilbao-Ubillos, 2008; Tian et al., 2024). Extensive research has substantiated that a multitude of maneuvers can instigate traffic oscillations, including lane changing (Ahn and Cassidy, 2007; Mauch and J. Cassidy, 2002; Laval and Daganzo, 2006), merging, and diverging (Ahn et al., 2010; Cassidy and Rudjanakanoknad, 2005). The rapid progression of vehicle automation technologies, particularly the advent of automated vehicles (AVs) (Guanetti et al., 2018; Rajamani, 2011), introduces a new traffic flow system: mixed traffic comprising a heterogeneous mix of AVs and human-driven vehicles (HDVs). This mixed traffic system is expected to reshape traditional traffic flow dynamics in unparalleled ways and improve traffic throughput, safety as well as energy efficiency during traffic oscillations (Talebpour and Mahmassani 2016; Li et al. 2024a; Xu et al. 2022). Consequently, acquiring a profound understanding of the distinctive behaviors exhibited by AVs during traffic oscillations becomes imperative in order to model and analyze this new traffic flow system.
@@
The aforementioned and similar work relied on linear models, which provide an elegant way of designing string-stable controllers in scenarios where the saturation limits on traffic state and control input (i.e., velocity and acceleration) are not reached. However, from an analysis perspective, such methods might produce unrealistic results. Consider a simple extreme example: suppose a vehicular string is string-unstable such that, at a specific frequency, the oscillation amplification ratio of each vehicle is greater than one, then as the number of vehicles increases, the amplitude of oscillations would approach infinity at the end of the vehicular string, making the acceleration/deceleration exceed the reasonable limit and vehicle may travel backward or surpass free flow speed. This highlights the necessity of analyzing the oscillatory response of AVs considering the presence of control and traffic state saturation limits, which are inherently nonlinear. Such limits are inherent in vehicular systems due to the physical limits on acceleration, deceleration, and velocity (Stone and Ball, 2004). Other nontrivial sources of saturation arise from the clipping of control signals generated by an unconstrained controller and the imposition of control and/or state constraints by constrained optimal controllers. These saturations are typically imposed to accommodate the aforementioned physical limits or to curb overly aggressive behaviors of AVs (Li et al., 2024b; Wang et al., 2014). It is crucial to note that saturation limits can be reached not only in string-unstable platoons, where oscillations are amplified along the vehicular string, but also in string-stable, heterogeneous platoons (Naus et al., 2010). particularly when an AV (e.g., automated truck) has more restrictive limits compared to its predecessor. For example, take an automated truck with a maximum absolute acceleration of $1\,\text{m/s}^{2}$ that follows a passenger car, which freely oscillates with a maximum absolute acceleration of $3\,\text{m/s}^{2}$. A string stable controller could generate an acceleration command ranging from 2-2 to $2\,\text{m/s}^{2}$, yet this still surpasses the acceleration limits of the automated truck.
@@ 2 Modeling the CF problem of AVs
[…]
@@
While Eq. (7) has also been applied to describe HDVs in the literature, they typically rely on strong assumptions about deterministic behavior and further linearization simplifications. To maintain a clear and rigorous scope, this study focuses on AVs.
@@
In practice, time delays could impact the response of AV systems. As this paper does not address connectivity issues, focusing on actuation delays is sufficient. We can introduce an actuation delay into our framework via the block diagram shown in Fig. 2, where the added delay block signifies an actuation delay of $\tau$ seconds. A time delay is known to be linear time-invariant, with a frequency response magnitude of 1 and a phase shift of $\omega\tau$ (Franklin et al., 2002). Actuation delay can be effectively compensated in lower-level control design (Xing et al., 2018; Xu et al., 2020), making it negligible for response analysis. Therefore, to simplify the presentation of our framework, we do not incorporate the delay component in our analysis that follows. It is worth noting that, the steps of our analysis procedure remain consistent even when incorporating the delay, as it merely represents an additional linear component without introducing any further nonlinearity.
@@ 3.2.1 Control saturation
For the control saturation element (Saturation 1 in Fig. 2), let the input of it be denoted as $a_{n+1}(t)=B_{a}sin(\omega t)$, considering the acceleration limits as $a_{min},a_{max}]$. The DF of the control saturation, $N_{a,sat}(B_{a})$, can be characterized into four cases corresponding to different limit activeness, depending on the value of $B_{a}$, given the known values of $a_{min}$ and $a_{max}$:
@@
Case (2): limit activeness by $a_{min}$ (i.e., $a_{min}\leq B_{a}\leq a_{max}$), as illustrated in Fig. 3.2.1(b), where $\beta_{1}=sin^{-1}(\frac{-a_{min}}{B_{a}})$.
@@
[…]
@@
The cases in this subsection only describe how a signal passes through the saturation element without considering the closed-loop effect. The closed-loop response is inherently much more complex than the open-loop case due to feedback interactions (illustrated in Fig. 2). A full characterization of the closed-loop behavior requires analyzing how signals propagate throughout the feedback loop. This is systematically addressed in Section 4, where we examine the overall system response.
@@ 3.2.2 State saturation
Similar, the DF of the state saturation (Saturation 2 in Fig. 2) with input $\tilde{v}_{n+1}(t)=B_{v}\sin(\omega t)$ and its output limited within $\tilde{v}_{min},\tilde{v}_{max}]$ can be derived in the same manner:
@@
Case (2): limit activeness by $\tilde{v}_{min}$ (i.e., $\tilde{v}_{min}\leq B_{v}\leq\tilde{v}_{max}$).
@@ 4.2.1 Systems with solely control saturation
$=\sqrt{(k_{1}R)^{2}+(\omega k_{2}R)^{2}}\sin\left(\omega t+\tan^{-1}\left(\frac{k_{2}\omega}{k_{1}}\right)\right)\qquad\qquad\text{if }1-\frac{k_{1}|N_{a,sat}(B_{a})|}{\omega^{2}}\geq 0$
@@
$=\sqrt{(k_{1}R)^{2}+(\omega k_{2}R)^{2}}\sin\left(\omega t+\tan^{-1}\left(\frac{k_{2}\omega}{k_{1}}\right)\right)\qquad\qquad\text{if }1-\frac{k_{1}|N_{a,sat}(B_{a})|}{\omega^{2}}<0$
@@ 4.2.2 Systems with solely state saturation
$=\sqrt{\left(\frac{-k_{1}R}{\omega}\right)^{2}+(k_{2}R)^{2}}\sin\left(\omega t-\tan^{-1}\left(\frac{k_{1}}{k_{2}\omega}\right)\right)\qquad\qquad\text{if }1-\frac{k_{1}|N_{v,sat}(B_{v})|}{\omega^{2}}\geq 0$
@@
$=\sqrt{\left(\frac{-k_{1}R}{\omega}\right)^{2}+(k_{2}R)^{2}}\sin\left(\omega t-\tan^{-1}\left(\frac{k_{1}}{k_{2}\omega}\right)\right)\qquad\qquad\text{if }1-\frac{k_{1}|N_{v,sat}(B_{v})|}{\omega^{2}}<0$
@@ 4.2.3 Systems with both control and state saturation
$\cdot\sin\left(\omega t+\phi_{a}-\tan^{-1}\left(\frac{-\omega k_{3}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}{\omega^{2}-k_{1}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}\right)\right)\qquad\qquad\text{if }1-\frac{k_{1}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}{\omega^{2}}\geq 0$
@@
$\cdot\sin\left(\omega t+\phi_{a}-\tan^{-1}\left(\frac{-\omega k_{3}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}{\omega^{2}-k_{1}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}\right)+\pi\right)\qquad\qquad\text{if }1-\frac{k_{1}|N_{v,sat}(B_{v})||N_{a,sat}(B_{a})|}{\omega^{2}}<0$
@@
$\omega^{2}\tilde{p}_{n+1}(\omega)=k_{1}\tilde{p}_{n}(\omega)+j\omega k_{2}\tilde{p}_{n}(\omega)-k_{1}\tilde{p}_{n+1}(\omega)+j\omega k_{3}\tilde{p}_{n+1}(\omega)$ (52)
@@ 4.2.4 Solving for forced oscillation candidates
For the three types of systems analyzed above, given the values of $B_{a,s}$ or $B_{v,s}$, the variables $\phi_{a,s}$ or $\phi_{v,s}$ can be analytically determined with ease by equating the phase of both sides of equations Eqs. (35)-(36), Eqs. (41)-(42), or Eqs. (47)-(48) within the studied range of $\omega$. $B_{a,s}$ or $B_{v,s}$ can be deduced by equating the amplitude of the two sides of the aforementioned equations. However, the nonlinearity introduced by the DFs can render the analytical solution of these equations challenging. Several studies have proposed graphical methods to identify the roots of such equations (Vander Velde et al., 1968; Atherton, 2011; Li et al., 2012; Maraini and Nataraj, 2018). Nevertheless, implementing these graphical solutions for heterogeneous AVs can be cumbersome and repetitive. Fortunately, iterative strategies, such as the Brent-Dekker method (Dekker, 1969; Brent, 2013), offer numerical solutions. These strategies are largely dependent on an initial guess, hence, we sweep the initial guess across the range $0,B_{ini,max}]$ to identify all possible solutions, where $B_{ini,max}$ represents a predetermined maximum value of the initial guess.
@@ 5.1 The incremental-input DF
where the first term is the predominant input which is one of the oscillation candidates acquired from the analysis in Section 4. It’s shifted on the time axis so that its first cycle starts at $t=0$, simplifying notation without affecting the analysis due to our interest in the phase relation between primary and incremental inputs. The second term represents the incremental input which can be viewed as a perturbation to the predominant input. In this context, $\epsilon\ll B_{s}$, indicating the perturbation’s magnitude is notably less than the predominant input. Additionally, $\theta$ represents the phase shift of the incremental input in relation to the predominant input, which can span any value in the interval $0,2\pi]$.
Remark 10.
Due to the low-pass characteristic of AV systems and the assumption of no subharmonics on the saturation, the output of the saturation remains sinusoidal with the same frequency as the external input to the AV system (i.e., the oscillatory position of the AV’s predecessor). The input to the nonlinear element, being a linear combination of the external input and feedback signals, also remains sinusoidal at the same frequency by trigonometric identities. Therefore, the predominant and perturbation terms are both considered sinusoidal with the same frequency, and perturbations are analyzed with all possible phases.
@@ 5.2.1 Control saturation
For the control saturation element, let the predominant input of it be denoted as $a_{n+1}(t)=B_{a,s}sin(\omega t)$, where $B_{a,s}$ can be derived from Eqs. (35)-(36) or Eqs. (47)-(48). Considering the acceleration limits as $a_{min},a_{max}]$. It has been shown that the DF of the control saturation, $N_{a,sat}(B_{a,s})$, can be characterized into four cases depending on different limit activeness. Analogously, the incremental-input DF, $N_{inc}(B_{a,s},\theta_{a})$, can also be classified into these four scenarios:
@@
Case (2): limit activeness by $a_{min}$ (i.e., $a_{min}\leq B_{a,s}\leq a_{max}$).
@@ 5.2.2 State saturation
Similarly, consider the predominant input to the state saturation $\tilde{v}_{n+1}(t)=B_{v,s}\sin(\omega t)$, where $B_{v,s}$ can be obtained by solving Eqs. (41)-(42) or Eqs. (47)-(48), and the output of the saturation limited within $\tilde{v}_{min},\tilde{v}_{max}]$. The incremental-input DF of the state saturation, $N_{v,inc}(B_{v,s},\theta_{v})$, can be obtained as follows:
@@
Case (2): limit activeness by $\tilde{v}_{min}$ (i.e., $\tilde{v}_{min}\leq B_{v,s}\leq\tilde{v}_{max}$).
@@ 6.1 Simulation of system containing one saturation
The results demonstrate that upon reaching the saturation limits, the linear theoretical method, which is commonly employed in much of the current research on string stability, significantly overestimates the magnitude of the frequency response, $|F|$. In other words, an underestimation of the attenuation effect in traffic oscillation largely exists for linear analysis. On the other hand, the DF approximation introduced by Zhou et al., 2023 marginally overestimates the oscillation attenuation. Both the linear method and DF approximation considerably overestimate the phase of the frequency response, $\angle F$. Consequently, the response time (defined as the duration before the AV responds to a predecessor’s driving behavior change), which can be calculated by $\frac{\angle F}{\omega}$, is underestimated.